Juniper Enterprises sells handmade clocks. Its variable cost per clock is $23.80, and each clock sells for $34.

 Calculate Juniper's contribution margin per unit and contribution margin ratio. If the company's fixed costs total $7,446, determine how many clocks Juniper must sell to break even 

Calculate Juniper's contribution margin per unit and contribution margin ratio. (Round your "Unit Contribution Margin" answer to 2 decimal places.) 


If the company's fixed costs total $7,446, determine how many docks Juniper must sell t break even.
Answer #1

Unit contribution margin

= Selling price per unit - Variable cost per unit

= 34 - 23.80

10.20

Contribution margin ratio

= Contribution margin/Selling price

= 10.20/34

30%

.

Breakeven units

= Fixed cost/Contribution margin per unit

= 7,446/10.20

730